Type your answer in the box. Suppose that nominal GDP last year was $25 million and is $27 million this year. Also suppose that inflation between last year and this year was 10%. This means that real GDP this year (using last years prices) is $ million (round to two decimal points).
Answer #1

Ans: $24.55 million

Explanation:

In the base year (here it is last year) NGDP is equal to RGDP

The GDP Deflator in last year = (NGDP / RGDP) * 100 = (25 million / 25 million) * 100 = 100

The inflation between last year and this year was 10% means, the GDP deflator in this year was 110

So, RGDP this year = (NGDP * 100) / GDP deflator = (27 * 100) / 110 = $24.55 million